OpenStudy (anonymous):

evaluate 8 (sigma symbol) n+25 n=3

OpenStudy (agent0smith):

\[\Large \sum_{3}^{8} n+25\] This means you want the sum of n+25, starting from n=3 and stopping at n=8: \[\Large \sum_{3}^{8} n+25 =\]\[ (3+25) + (4+25) + (5+25) + (6+25) + (7+25) + (8+25)=\]

OpenStudy (anonymous):

thankyou!!:)

OpenStudy (agent0smith):

Welcome :)